# Preference Tuning with Human Feedback on Language, Speech, and Vision Tasks: A Survey

### v1

@@ 7.2.1 VHELM
VHELM [^1] https://crfm.stanford.edu/helm/vhelm/latest/ is an extension of the HELM framework (Liang et al., 2022) with the adaptation methods to assess the performance of VLMs by scoring the winning rates against the GPT-4V model.
@@ 7.4 Reward Model Evaluation
One way to assess the quality of our model is by evaluating the performance of the reward model using a benchmark. Zhu et al. (2024); Jiang et al. (2023b) propose using validation sets from previous RLHF training processes, such as Anthropic’s Helpful and Harmless data (Bai et al., 2022a) or OpenAI’s Learning to Summarize (Stiennon et al., 2020). Additionally, newly released preference data, aimed at expanding the diversity of preference training datasets, such as UltraFeedback (Cui et al., 2023), UltraInteract (Yuan et al., 2024a), and Nectar (Zhu et al., 2024), lack test sets, necessitating a new style of evaluation for reward models. RewardBench is a benchmark dataset and codebase designed for this purpose (Lambert et al., 2024). The dataset comprises a collection of prompt-chosen-rejected triplets that span various domains, including chat, reasoning, and safety. This allows for a comprehensive evaluation of how reward models perform on challenging, structured, and out-of-distribution queries.
@@ 8.2 Research Directions
8.2.1 Multilinguality
While significant resources have been allocated to enhance the safety of LLMs for deployment, the safety of multilingual LLMs remains underexplored. Ahmadian et al. (2024b) is one of the pioneering works pushing the boundaries of aligning language models by optimizing for both general and safety performance simultaneously in a multilingual setting using Distributional DPO. Similarly, Li et al. (2024) propose exploring DPO training to reduce toxicity in multilingual open-ended generations. Another line of research focuses on using multilingual alignment based on human preferences to improve reasoning abilities, aiming to align reasoning processes in other languages with those in the dominant language (She et al., 2024). There is still ample room for exploration in the multilingual space, particularly in examining the cultural aspects of multilingualism (Adilazuarda et al., 2024; AlKhamissi et al., 2024) and improving the alignment of LLM for generation (Winata et al., 2021b). It is crucial to cover more diverse languages, including regional languages, different dialects (Aji et al., 2022), and code-switching (Winata et al., 2021a), which are common phenomena in bilingual and multilingual communities (Winata et al., 2024). Additionally, the exploration of multilingual topics in vision-language and speech tasks remains open for further investigation.
@@ 8.2.6 Mechanistic Understanding of Preference Tuning Methods
Acknowledgement: Wenpin Tang and Hanyang Zhao are supported by NSF grant DMS-2206038, a start-up grant at Columbia University, and the Columbia Innovation Hub grant. The works of Hanyang Zhao and David D. Yao are part of a Columbia-CityU/HK collaborative project that is supported by InnotHK Initiative, The Government of the HKSAR and the AIFT Lab.

### v2

@@ 7.2.1 VHELM
VHELM [^1] https://crfm.stanford.edu/helm/vhelm/latest/. is an extension of the HELM framework (Liang et al., 2022) with the adaptation methods to assess the performance of VLMs by scoring the winning rates against the GPT-4V model.
@@ 7.4 Reward Model Evaluation
One way to assess the quality of our model is by evaluating the performance of the reward model using a benchmark. Zhu et al. (2024); Jiang et al. (2023b) propose using validation sets from previous RLHF training processes, such as Anthropic’s Helpful and Harmless data (Bai et al., 2022a) or OpenAI’s Learning to Summarize (Stiennon et al., 2020). Additionally, newly released preference data, aimed at expanding the diversity of preference training datasets, such as UltraFeedback (Cui et al., 2023), UltraInteract (Yuan et al., 2024a), and Nectar (Zhu et al., 2024), lack test sets, necessitating a new style of evaluation for reward models. RewardBench is a benchmark dataset and codebase designed for this purpose (Lambert et al., 2024). The dataset comprises a collection of prompt-chosen-rejected triplets that span various domains, including chat, reasoning, and safety. This allows for a comprehensive evaluation of how reward models perform on challenging, structured, and out-of-distribution queries. Winata et al. (2024a) propose MetaMetrics, a new method to construct a meta-metric that is aligned with human preferences by calibrating multiple metrics by using Bayesian optimization and boosting methods, which has been further applied to machine translation (Anugraha et al., 2024).
@@ 8.2 Research Directions
8.2.1 Multilingual, Multicultural, and Pluralistic Preference Tuning
While significant resources have been allocated to enhance the safety of LLMs for deployment, the safety of multilingual LLMs remains underexplored. Ahmadian et al. (2024b) is one of the pioneering works pushing the boundaries of aligning language models by optimizing for both general and safety performance simultaneously in a multilingual setting using Distributional DPO. Similarly, Li et al. (2024) propose exploring DPO training to reduce toxicity in multilingual open-ended generations. Another line of research focuses on using multilingual alignment based on human preferences to improve reasoning abilities, aiming to align reasoning processes in other languages with those in the dominant language (She et al., 2024). There is still ample room for exploration in the multilingual space, particularly in examining the cultural aspects of multilingualism (Adilazuarda et al., 2024; AlKhamissi et al., 2024) and improving the alignment of LLM for generation (Winata et al., 2021b). It is crucial to cover more diverse languages, including regional languages, different dialects (Aji et al., 2022), and code-switching (Winata et al., 2021a), which are common phenomena in bilingual and multilingual communities (Winata et al., 2024b). Additionally, the exploration of multilingual topics in vision-language and speech tasks remains open for further investigation.
@@ 8.2.6 Mechanistic Understanding of Preference Tuning Methods
Acknowledgments: Wenpin Tang and Hanyang Zhao are supported by NSF grant DMS-2206038, a start-up grant at Columbia University, and the Columbia Innovation Hub grant. The works of Hanyang Zhao and David D. Yao are part of a Columbia-CityU/HK collaborative project that is supported by InnotHK Initiative, The Government of the HKSAR and the AIFT Lab.